# An Empirical Study of Mamba-based Pedestrian Attribute Recognition

### v1

@@ Abstract
Current strong pedestrian attribute recognition models are developed based on Transformer networks, which are computationally heavy. Recently proposed models with linear complexity (e.g., Mamba) have garnered significant attention and have achieved a good balance between accuracy and computational cost across a variety of visual tasks. Relevant review articles also suggest that while these models can perform well on some pedestrian attribute recognition datasets, they are generally weaker than the corresponding Transformer models. To further tap into the potential of the novel Mamba architecture for PAR tasks, this paper designs and adapts Mamba into two typical PAR frameworks, i.e., the text-image fusion approach and pure vision Mamba multi-label recognition framework. It is found that interacting with attribute tags as additional input does not always lead to an improvement, specifically, Vim can be enhanced, but VMamba cannot. This paper further designs various hybrid Mamba-Transformer variants and conducts thorough experimental validations. These experimental results indicate that simply enhancing Mamba with a Transformer does not always lead to performance improvements but yields better results under certain settings. We hope this empirical study can further inspire research in Mamba for PAR, and even extend into the domain of multi-label recognition, through the design of these network structures and comprehensive experimentation. The source code of this work will be released at https://github.com/Event-AHU/OpenPAR
1 Introduction
Pedestrian Attribute Recognition (PAR) [48] is a widely exploited research topic in the computer vision (CV) community. It aims to recognize human attributes from a set of attribute descriptions, such as short black hair, wearing hats, with back bag, etc. On the one hand, pedestrian attribute recognition can be used to describe the appearance and motion characteristics of pedestrians, playing a crucial role in understanding them; on the other hand, pedestrian attributes can serve as mid-level semantic representations, assisting in improving the performance of other visual tasks, including pedestrian detection [55], person re-identification [54, 59], and more.
With the help of deep learning, various deep PAR models are proposed, including CNN (Convolutional Neural Networks), RNN (Recurrent Neural Networks), GNN (Graph Neural Networks), Transformer, etc. Specifically, MTCNN [1] uses CNN to extract features and combine them in a multi-task manner to solve the PAR task. JRL [44] and GRL [56] are solving PAR as a sequence prediction task using LSTM to model the association between attributes. PARformer [7] and VTB [3] use Transformer to solve PAR task, the difference is that VTB introduces text modality. Although existing PAR models work well in simple scenarios, the recognition performance in challenging scenarios is still limited (e.g., low illumination, cross-domain). In addition, the training/inference cost ($\mathcal{O}(N^{2})$) is rather high due to the utilization of self-attention in Transformer networks. Based on the observations above and reflections, it is natural to raise the following question: how can we design a new pedestrian attribute recognition framework that achieves a comparable or even better performance compared to the Transformer-based models, meanwhile, reducing the cost significantly?
Recently, the State Space Model (SSM) has drawn more and more attention in the artificial intelligence community due to its linear complexity ($\mathcal{O}(N)$) and good performance. It has been widely utilized for visual tracking [17], image-/video-based classification [60, 27], and time series analysis [49]. The experimental results reported in the SSM survey [47] demonstrate that the Vim-S [60] based attribute recognition model performs better than the ViT-S based version on the PA100K [31] dataset, but achieves worse results on the PETA [5] dataset. Hence, crafting innovative Mamba-inspired PAR architectures that deliver uniform performance enhancements across benchmarks, while also maintaining low training and testing overhead, poses an ongoing challenge.
@@ 2 Related Works
2.1 Pedestrian Attribute Recognition
Pedestrian attribute recognition has made significant progress over the years, with many methods proposed by researchers achieving promising results. Early methods primarily involved feeding images into CNN networks for multi-task training, and sharing parameters between networks to obtain richer features. For example, MTCNN [1] adopted this multi-task training approach, using CNN to extract image features and allowing knowledge sharing among different attribute categories. Due to the correlations between different attributes, sequential prediction models were introduced into pedestrian attribute recognition research. JRL [44] modeled the contextual information between pedestrians and the contextual information between attributes using LSTM, transforming pedestrian attribute recognition into a sequence prediction problem. Building upon JRL, GRL [56] divided the attribute list into multiple attribute groups and modeled the spatial and semantic correlations between groups using LSTM. With the development of deep learning networks, attention models began to be used in pedestrian attribute recognition. Specifically, HydraPlus-Net [32] introduced a multi-directional attention module to extract multi-scale attention features for obtaining more comprehensive pedestrian features. Transformer, a model proposed by Vaswani et al. [43], has been widely used in the field of natural language processing and has achieved great success. Some researchers have also applied it to pedestrian attribute recognition. Fan et al. [7] proposed a pure Transformer-based PAR network, using a Transformer-based backbone network for image feature extraction to obtain more discriminative features. Cheng et al. [3] proposed VTB, encoding pedestrian images into visual features and encoding attribute annotations into text features using pre-trained text encoders, and using Transformer for multimodal fusion. However, due to the global attention mechanism of the Transformer, its complexity grows quadratically with the length of the input sequence. Gu et al. [9] introduced Mamba, a type of State Space Model that optimizes the complexity of sequence processing to linear. Therefore, we introduce a state space modal for optimization, improving computational efficiency while ensuring the accuracy of pedestrian attribute recognition.
2.2 State Space Model
The State Space Model (SSMs) [13, 12, 11, 9] is a mathematical framework used to model dynamic systems, which has received widespread attention in recent years due to its excellent performance in handling long sequence data. S4 [12], as a pioneering work in using state space modal for long sequence processing, introduced Hippo [10] to address the issue of long-range dependencies in sequence modeling, significantly improving model performance. S4ND [34] extends the spatial state model to multidimensional signals, enabling it to model large-scale visual data as continuous multidimensional signals, laying the foundation for SSMs’s use in multimodal tasks. Mamba [9] introduces a selective mechanism for information processing while balancing training, inference efficiency, and model effectiveness, outperforming Transformer [43] in natural language processing. After Mamba, there appeared works applying Mamba to the visual domain, such as Vim [60] and VMamba [33]. Both of these works show the potential of Mamba in vision tasks. Vim proposes a bidirectional Mamba to model images, and VMamba proposes a four-direction cross-scan module. Through this multi-direction sequence order, it is more effective to model two-dimensional image data and maintain linear computational complexity. Mamba-2 [4] is the improvement of Mamba, and proposes SSD (State Space Duality) framework, which improves Mamba by 2-8 times speedup and obtains better performance. Our work explores the potential of Mamba in handling visual-text fusion features and comprehensively applies Mamba and its derivative modules to pedestrian attribute recognition.
3 Methodology
@@
3.1 Preliminary: State Space Model
@@
3.2 When Mamba Meets PAR
@@
3.3 Hybrid Mamba-Transformer for PAR
In our experiments, we find that the pure VMamba [33] based PAR framework performs better than the image-text fusion based one. Therefore, in this section, we exploit new network architectures to aggregate the VMamba [33] with Transformer networks for better pedestrian attribute recognition. Specifically speaking, we design eight variations of hybrid Mamba-Transformer networks to validate which one performs better for the PAR task. Due to the limited space in this paper, we briefly introduced the computation procedures of these hybrid Mamba-Transformer networks for the PAR task. For more details on these models, please check our source code.
@@ 3.3 Hybrid Mamba-Transformer for PAR
4 Experiments
4.1 Datasets and Evaluation Metric
In our experiments, seven widely used PAR benchmark datasets are evaluated, including PA100K [31], PETA [5], RAP-V1 [26], RAP-V2 [25], WIDER [29], PETA-ZS [20], and RAP-ZS [20]. A brief introduction to these datasets is given below:
@@ 4.2 Implementation Details
For the detailed parameters, we train the model for 100 epochs using the Adam [37] optimizer. When using VMamba-B, we set the learning rate to 8e-5 and the batch size to 16. When using Vim-S, we set the learning to 2.5e-5 and the batch size to 64. Our source code is implemented using Python and the deep learning framework PyTorch [35]. The experiments were conducted on a server with GPU RTX3090s. More details can be found in our source code.
4.3 Comparison on Public PAR Benchmarks
@@
Table 1: Comparison with state-of-the-art methods on PETA and PA100K datasets. “V” means Vision branch, “L” means Language branch. “-” means this indicator is not available.
Methods Backbone PETA PA100K
mA Acc Prec Recall F1 mA Acc Prec Recall F1
DeepMAR (ACPR 2015) [24] CaffeNet 82.89 75.07 83.68 83.14 83.41 72.70 70.39 82.24 80.42 81.32
HPNet (ICCV 2017) [31] Inception 81.77 76.13 84.92 83.24 84.07 74.21 72.19 82.97 82.09 82.53
JRL (ICCV 2017) [44] AlexNet 82.13 - 82.55 82.12 82.02 - - - - -
GRL (IJCAI 2018) [56] Inception-V3 86.70 - 84.34 88.82 86.51 - - - - -
MsVAA (ECCV 2018) [38] ResNet101 84.59 78.56 86.79 86.12 86.46 - - - - -
RA (AAAI 2019) [57] Inception-V3 86.11 - 84.69 88.51 86.56 - - - - -
VRKD (IJCAI 2019) [28] ResNet50 84.90 80.95 88.37 87.47 87.91 77.87 78.49 88.42 86.08 87.24
AAP (IJCAI 2019) [16] ResNet50 86.97 79.95 87.58 87.73 87.65 80.56 78.30 89.49 84.36 86.85
VAC (CVPR 2019) [15] ResNet50 - - - - - 79.16 79.44 88.97 86.26 87.59
ALM (ICCV 2019) [41] BN-Inception 86.30 79.52 85.65 88.09 86.85 80.68 77.08 84.24 88.84 86.46
JLAC (AAAI 2020) [40] ResNet50 86.96 80.38 87.81 87.09 87.50 82.31 79.47 87.45 87.77 87.61
SCRL (TCSVT 2020) [51] ResNet50 87.2 - 89.20 87.5 88.3 80.6 - 88.7 84.9 82.1
SSCsoft (ICCV 2021) [18] ResNet50 86.52 78.95 86.02 87.12 86.99 81.87 78.89 85.98 89.10 86.87
IAA-Caps (PR 2022) [50] OSNet 85.27 78.04 86.08 85.80 85.64 81.94 80.31 88.36 88.01 87.80
MCFL (NCA 2022) [2] ResNet-50 86.83 78.89 84.57 88.84 86.65 81.53 77.80 85.11 88.20 86.62
DRFormer (NC 2022) [42] ViT-B/16 89.96 81.30 85.68 91.08 88.30 82.47 80.27 87.60 88.49 88.04
VAC-Combine (IJCV 2022) [14] ResNet50 - - - - - 82.19 80.66 88.72 88.10 88.41
DAFL (AAAI 2022) [19] ResNet50 87.07 78.88 85.78 87.03 86.40 83.54 80.13 87.01 89.19 88.09
CGCN (TMM 2022) [6] ResNet 87.08 79.30 83.97 89.38 86.59 - - - - -
CAS-SAL-FR (IJCV 2022) [53] ResNet50 86.40 79.93 87.03 87.33 87.18 82.86 79.64 86.81 87.79 85.18
PARformer (TCSVT 2023) [7] Swin-L 89.32 82.86 88.06 91.98 89.06 84.46 81.13 88.09 91.67 88.52
VTB (TCSVT 2022) [3] ViT-B/16 85.31 79.60 86.76 87.17 86.71 83.72 80.89 87.88 89.30 88.21
VTB* (TCSVT 2022) [3] ViT-L/14 86.34 79.59 86.66 87.82 86.97 85.30 81.76 87.87 90.67 88.86
SequencePAR (arXiv-2023) [21] ViT-L/14 - 84.92 90.44 90.73 90.46 - 83.94 90.38 90.23 90.10
Results on PETA dataset. As shown in Table 1, the evaluated MambaPAR model performs well on most evaluation metrics, with different branches and backbone networks showing varying results. Specifically, the VMamba-B visual branch model achieved mA, Acc, Prec, Recall, and F1 metrics of 86.28, 80.54, 87.45, 87.95, and 87.45, respectively; the VMamba-B visual plus language branch model achieved 85.01, 78.47, 85.12, 87.49, and 86.00; the Vim visual plus language branch model achieved 81.45, 74.92, 82.68, 84.27, and 83.15; and the Vim visual plus text branch model achieved 84.25, 76.07, 82.73, 87.20, and 84.56. Interestingly, we can find that the attribute labels are useful for the Vim-S based PAR, but not for the VMamba-B based version. Compared to the Transformer-based pedestrian attribute recognition models, our VMamba visual branch model surpassed VTB (ViT-B/16, 85.31/79.60/86.76/87.17/86.71) and VTB* (ViT-L/14, 86.34/79.59/86.66/87.82/86.97) in the mA, Acc, Prec, Recall, and F1 metrics. Therefore, we can conclude that our model achieves competitive results on the PETA dataset, verifying the feasibility of the Mamba model for pedestrian attribute recognition.
Results on PA100K dataset. As shown in Table 1, our MambaPAR model also performs well on the PA100K dataset. The VMamba-B visual branch model achieved mA, Acc, Prec, Recall, and F1 metrics of 83.63, 81.11, 87.59, 89.9, and 88.40, respectively; the VMamba-B visual plus language branch model achieved 81.50, 79.55, 87.54, 88.03, and 87.35; the Vim visual plus language branch model achieved 79.28, 76.77, 85.24, 86.58, and 85.44; and the Vim visual plus language model achieved 80.87, 79.33, 86.48, 88.79, and 87.21. The VMamba-B visual branch model’s performance is close to that of VTB* (85.3/81.76/87.87/90.67/88.86). These experimental results indicate that the exploited Mamba-based PAR frameworks we explored have achieved preliminary success.
From the experimental results reported in the PETA and PA100K datasets, we can find that the VMamba-B achieves the best results over other Mamba-based PAR models. Therefore, we only report the results of pure VMamba-B based PAR framework for the rest of the datasets.
Results on RAP-V1 dataset. As shown in Table 2, our VMamba-B visual branch model achieved mA, Acc, Prec, Recall, and F1 metrics of 83.12, 69.47, 78.03, 84.81, and 80.88, respectively. In comparison, the strong baseline VTB* achieved 83.69, 69.78, 78.09, 85.21, and 81.10 on these metrics, and our results are very close to theirs. Meanwhile, our model’s performance is also close to the DRFormer [42], which is also developed based on Transformer. This fully demonstrates that the performance of our Mamba-based pedestrian attribute recognition model is comparable to or even surpasses Transformer-based models.
Results on RAP-V2 dataset. As shown in Table 2, our VMamba-B visual branch model achieved mA, Acc, Prec, Recall, and F1 metrics of 81.91, 68.08, 76.41, 84.38, and 79.83, respectively, outperforming the Transformer-based VTB* (81.36/67.58/76.19/84.00/79.52). This further proves the effectiveness of our proposed pedestrian attribute recognition model.
Table 2: Comparison with state-of-the-art methods on RAPv1 and RAPv2 datasets. “V” means Vision branch. “-” means this indicator is not available.

### v2

@@ Abstract
Current strong pedestrian attribute recognition models are developed based on Transformer networks, which are computationally heavy. Recently proposed models with linear complexity (e.g., Mamba) have garnered significant attention and have achieved a good balance between accuracy and computational cost across a variety of visual tasks. Relevant review articles also suggest that while these models can perform well on some pedestrian attribute recognition datasets, they are generally weaker than the corresponding Transformer models. To further tap into the potential of the novel Mamba architecture for PAR tasks, this paper designs and adapts Mamba into two typical PAR frameworks, i.e., the text-image fusion approach and pure vision Mamba multi-label recognition framework. It is found that interacting with attribute tags as additional input does not always lead to an improvement, specifically, Vim can be enhanced, but VMamba cannot. This paper further designs various hybrid Mamba-Transformer variants and conducts thorough experimental validations. These experimental results indicate that simply enhancing Mamba with a Transformer does not always lead to performance improvements but yields better results under certain settings. We hope this empirical study can further inspire research in Mamba for PAR, and even extend into the domain of multi-label recognition, through the design of these network structures and comprehensive experimentation. The source code and pre-trained models will be released on https://github.com/Event-AHU/OpenPAR.
Index Terms:
Mamba, PAR, Hybrid architecture
I Introduction
Pedestrian Attribute Recognition (PAR), as surveyed in [1], is a prevalent and significant research area within the computer vision (CV) community. The primary objective of PAR is to discern human attributes based on a variety of descriptive cues, such as short black hair, wearing hats, carrying a backpack, and so forth. This capability is twofold: firstly, it provides a detailed description of pedestrians’ appearances and movements, which is essential for a comprehensive understanding of their behavior; secondly, these attributes act as mid-level semantic representations that can significantly enhance the performance of various visual tasks. Such tasks include, but are not limited to, pedestrian detection [2], person re-identification [3, 4], and others.
Leveraging the power of deep learning, a multitude of sophisticated Pedestrian Attribute Recognition (PAR) models have been introduced, encompassing a range of architectures such as Convolutional Neural Networks (CNNs), Recurrent Neural Networks (RNNs), Graph Neural Networks (GNNs), and Transformers. For instance, the Multi-Task Convolutional Neural Network (MTCNN) [5] employs CNNs to extract features and integrates them in a multi-task learning framework to address the PAR challenge. Joint Representation Learning (JRL) [6] and Graph Representation Learning (GRL) [7] approach PAR as a sequence prediction problem, utilizing Long Short-Term Memory (LSTM) networks to capture the inter-attribute relationships. More recently, models like PARformer [8] and Visual-Textual Binding (VTB) [9] have adopted the Transformer architecture for PAR, with VTB notably incorporating textual information to enhance performance.
Despite the success of these existing models in straightforward scenarios, their effectiveness is often constrained in complex environments, such as those with poor lighting or cross-domain variations. Furthermore, the computational expense during training and inference, which scales with $\mathcal{O}(N^{2})$ due to the self-attention mechanism inherent in Transformer models, remains a significant challenge.
Given these considerations, it is pertinent to inquire: How can we devise a novel framework for pedestrian attribute recognition that not only rivals or surpasses the performance of Transformer-based models but also substantially reduces the associated computational costs? This question drives the pursuit of innovative solutions that balance efficiency with accuracy in the realm of PAR.
The State Space Model (SSM) has recently garnered significant attention within the artificial intelligence community, particularly for its linear complexity ($\mathcal{O}(N)$) and impressive performance across various applications. It has been extensively applied to tasks such as visual tracking [10], image and video-based classification [11, 12], and time series analysis [13]. The SSM survey [14] presents compelling evidence that models like Vim-S [11], which leverage SSM for attribute recognition, outperform their ViT-S counterparts on the PA100K [15] dataset, although they may underperform on the PETA [16] dataset. This disparity highlights the need for innovative architectures inspired by Mamba that can consistently enhance performance across different benchmarks while maintaining low computational costs for both training and testing. The development of such frameworks remains a challenging and active area of research, with the potential to significantly advance the field of pedestrian attribute recognition.
@@ II Related Works
II-A Pedestrian Attribute Recognition
Pedestrian attribute recognition has experienced remarkable advancements over the years, with a multitude of innovative methods proposed by the research community that have yielded encouraging outcomes. Initially, the primary approach involved the utilization of Convolutional Neural Networks (CNNs) for multi-task training, which facilitated the sharing of parameters across networks to enrich feature extraction. A prime example of this approach is MTCNN [5], which embraced multi-task training to extract image features using CNNs and fostered knowledge sharing across various attribute categories. Given the inherent correlations between different attributes, sequential prediction models were subsequently integrated into pedestrian attribute recognition research. JRL [6] employed Long Short-Term Memory (LSTM) networks to capture contextual relationships both between pedestrians and within attribute sets, effectively framing pedestrian attribute recognition as a sequence prediction challenge. Expanding on the foundation laid by JRL, GRL [7] segmented the attribute list into multiple groups and leveraged LSTM to model the spatial and semantic inter-group correlations.
As deep learning networks evolved, attention models started to be incorporated into pedestrian attribute recognition, leading to more nuanced feature extraction. Notably, HydraPlus-Net [Liu_2017_ICCV] incorporated a multi-directional attention module to distill multi-scale attention features, thereby acquiring a more comprehensive set of pedestrian attributes. The Transformer model, introduced by Vaswani et al. [23] and widely celebrated in the field of natural language processing, has also been successfully adapted for use in pedestrian attribute recognition. Fan et al. [8] proposed a purely Transformer-based network for PAR, employing a Transformer backbone for image feature extraction to obtain more discriminative features. Cheng et al. [9] introduced VTB, a model that encodes pedestrian images into visual features and attribute annotations into text features using pre-trained text encoders, subsequently employing Transformer for multimodal fusion.
However, the global attention mechanism characteristic of Transformers poses a challenge, as its computational complexity increases quadratically with the length of the input sequence. To address this, Gu et al. [24] introduced Mamba, a State Space Model designed to optimize sequence processing complexity to a linear scale. Consequently, we advocate for the adoption of a state space model to enhance optimization, thereby improving computational efficiency without compromising the accuracy of pedestrian attribute recognition.
II-B State Space Model
The State Space Model (SSMs) [25, 26, 27, 24] has emerged as a prominent mathematical framework for modeling dynamic systems, garnering significant attention for its exceptional capability to manage long sequence data. S4 [26], a groundbreaking approach in the application of state space models for long sequence processing, incorporated Hippo [28] to tackle long-range dependencies in sequence modeling, thereby markedly enhancing model performance. Building on this, S4ND [29] expanded the spatial state model to encompass multidimensional signals, allowing it to process large-scale visual data as continuous multidimensional signals. This development laid the groundwork for the application of SSMs in multimodal tasks.
Mamba [24] introduced a novel selective mechanism for information processing, striking a balance between training and inference efficiency while maintaining model effectiveness, and has been shown to surpass the Transformer [23] in natural language processing tasks. In the wake of Mamba’s success, several works have begun to apply its principles to the visual domain, such as Vim [11] and VMamba [17]. These efforts have demonstrated the potential of Mamba in vision-related tasks. Vim introduced a bidirectional Mamba model for image processing, while VMamba proposed a four-direction cross-scan module, which more effectively models two-dimensional image data through a multi-directional sequence order, all while preserving linear computational complexity.
Mamba-2 [30] represents an enhancement over the original Mamba model, introducing the SSD (State Space Duality) framework. This advancement has not only accelerated Mamba’s performance by 2-8 times but also achieved superior results.
Our research delves into the potential of Mamba for handling visual-text fusion features and applies Mamba and its derivative modules comprehensively to pedestrian attribute recognition. Through these explorations, we aim to leverage the strengths of Mamba in both visual and textual domains to enhance the accuracy and efficiency of attribute recognition in pedestrians.
III Our Proposed Approach
@@
III-A Preliminary: State Space Model
@@
III-B When Mamba Meets PAR
@@
III-C Hybrid Mamba-Transformer for PAR
In the experiments, we find the MambaPAR still has some performance differences directly with existing transformer-based methods. Therefore, in this section, we exploit the new network architecture to explore Mamba aggregation with the Transformer network for better pedestrian attribute recognition. Specifically speaking, we design eight variations of hybrid Mamba-Transformer networks to validate which one performs better for the PAR task. It should be noted that we do not use Vmamba in the hybrid architecture, but use Vim to mix with Transformer, first because the performance gap between Vim and Transformer is larger, and second because the GPU used in our experiments is a single RTX3090 with 24G memory. It cannot support our hybrid architecture experiment of VMamba and Transformer. Due to the limited space in this paper, we briefly introduced the computation procedures of these hybrid Mamba-Transformer networks for the PAR task. For more details on these models, please check our source code.
@@ III-C Hybrid Mamba-Transformer for PAR
TABLE I: Comparison with state-of-the-art methods on PETA and PA100K datasets. “V" means Vision branch, “L" means Language branch. “-" means this indicator is not available.
Methods Backbone PETA PA100K
mA Acc Prec Recall F1 mA Acc Prec Recall F1
DeepMAR [32] ACPR15 CaffeNet 82.89 75.07 83.68 83.14 83.41 72.70 70.39 82.24 80.42 81.32
HPNet [15] ICCV17 Inception 81.77 76.13 84.92 83.24 84.07 74.21 72.19 82.97 82.09 82.53
JRL [6] ICCV17 AlexNet 82.13 - 82.55 82.12 82.02 - - - - -
GRL [7] IJCAI18 Inception-V3 86.70 - 84.34 88.82 86.51 - - - - -
MsVAA [36] ECCV18 ResNet101 84.59 78.56 86.79 86.12 86.46 - - - - -
RA [37] AAAI19 Inception-V3 86.11 - 84.69 88.51 86.56 - - - - -
VRKD [38] IJCAI19 ResNet50 84.90 80.95 88.37 87.47 87.91 77.87 78.49 88.42 86.08 87.24
AAP [39] IJCAI19 ResNet50 86.97 79.95 87.58 87.73 87.65 80.56 78.30 89.49 84.36 86.85
VAC [40] CVPR19 ResNet50 - - - - - 79.16 79.44 88.97 86.26 87.59
ALM [41] ICCV19 BN-Inception 86.30 79.52 85.65 88.09 86.85 80.68 77.08 84.24 88.84 86.46
JLAC [42] AAAI20 ResNet50 86.96 80.38 87.81 87.09 87.50 82.31 79.47 87.45 87.77 87.61
SCRL [43] TCSVT20 ResNet50 87.2 - 89.20 87.5 88.3 80.6 - 88.7 84.9 82.1
SSCNet [44] ICCV21 ResNet50 86.52 78.95 86.02 87.12 86.99 81.87 78.89 85.98 89.10 86.87
IAA-Caps [45] PR22 OSNet 85.27 78.04 86.08 85.80 85.64 81.94 80.31 88.36 88.01 87.80
MCFL [46] NCA22 ResNet-50 86.83 78.89 84.57 88.84 86.65 81.53 77.80 85.11 88.20 86.62
DRFormer [33] NC22 ViT-B/16 89.96 81.30 85.68 91.08 88.30 82.47 80.27 87.60 88.49 88.04
VAC-Combine [47] IJCV22 ResNet50 - - - - - 82.19 80.66 88.72 88.10 88.41
DAFL [48] AAAI22 ResNet50 87.07 78.88 85.78 87.03 86.40 83.54 80.13 87.01 89.19 88.09
CGCN [49] TMM22 ResNet 87.08 79.30 83.97 89.38 86.59 - - - - -
CAS-SAL-FR [50] IJCV22 ResNet50 86.40 79.93 87.03 87.33 87.18 82.86 79.64 86.81 87.79 85.18
VTB [9] TCSVT22 ViT-B/16 85.31 79.60 86.76 87.17 86.71 83.72 80.89 87.88 89.30 88.21
VTB* [9] TCSVT22 ViT-L/14 86.34 79.59 86.66 87.82 86.97 85.30 81.76 87.87 90.67 88.86
PARformer [8] TCSVT23 Swin-L 89.32 82.86 88.06 91.98 89.06 84.46 81.13 88.09 91.67 88.52
OAGCN [51] TMM23 Swin-L 89.91 82.95 88.26 89.10 88.68 83.74 80.38 84.55 90.42 87.39
SequencePAR [52] arXiv23 ViT-L/14 - 84.92 90.44 90.73 90.46 - 83.94 90.38 90.23 90.10
PromptPAR [34] TCSVT24 ViT-L/14 88.76 82.84 89.04 89.74 89.18 87.47 83.78 89.27 91.70 90.15
SSPNet  [53] PR24 Swin-S 88.73 82.80 88.48 90.55 89.50 83.58 80.63 87.79 89.32 88.55
@@
IV Experiments
IV-A Datasets and Evaluation Metric
In our experiments, seven widely used PAR benchmark datasets are evaluated, including PA100K [15], PETA [16], RAP-V1 [18], RAP-V2 [19], WIDER [20], PETA-ZS [21], RAP-ZS [21], and MSP60K [54]. A brief introduction to these datasets is given below:
@@ IV-A Datasets and Evaluation Metric
TABLE II: Comparison with state-of-the-art methods on MSP60K datasets. "V" means Vision branch, "L" means Language branch.
Methods Backbone Random Split Cross-domain Split
mA Acc Prec Recall F1 mA Acc Prec Recall F1
DeepMAR [32] ACPR15 CaffeNet 70.46 72.83 84.71 81.46 83.06 54.84 44.97 63.38 58.81 61.01
RethinkingPAR [21] arXiv20 ResNet-101 74.01 74.20 84.17 83.94 84.06 55.98 46.52 62.85 62.09 62.47
SSCNet [44] ICCV21 ResNet-50 69.71 69.31 79.22 82.47 80.82 52.84 40.88 56.26 58.64 57.43
VTB [9] TCSVT22 ViT-B/16 76.09 75.36 83.56 86.46 84.56 58.59 49.81 65.11 66.11 65.00
DFDT [55] EAAI22 Swin-T 74.19 76.35 85.03 86.35 85.69 57.85 49.97 65.34 66.18 65.76
Zhou et al. [56] IJCAI23 ConvNeXt-B 73.07 68.76 78.38 82.10 80.20 54.26 41.91 56.23 60.11 58.11
PARformer [8] TCSVT23 Swin-L 76.14 76.67 84.77 86.93 85.44 57.96 50.63 62.28 71.04 65.82
SequencePAR [52] arXiv23 ViT-L/14 71.88 71.99 83.24 82.29 82.29 57.88 50.27 65.81 65.79 65.37
VTB-PLIP [57] arXiv23 ViT-B/16 73.90 73.16 82.01 84.82 82.93 56.30 46.77 61.20 64.47 62.18
Rethink-PLIP [57] arXiv23 ResNet-101 69.44 68.90 79.82 81.15 80.48 57.18 46.98 63.57 62.16 62.86
PromptPAR [34] TCSVT24 Vit-L/14 78.81 76.53 84.40 87.15 85.35 63.24 53.62 66.15 71.84 68.32
SSPNet  [53] PR24 Swin-S 74.03 74.10 84.01 84.02 84.02 56.15 46.75 62.44 63.07 62.75
MambaPAR (V+L) Vim-S 73.85 73.64 83.19 84.29 83.28 56.75 47.34 61.92 64.98 62.80
MambaPAR (V) VMamba-B 74.82 74.68 82.32 87.08 84.18 58.26 49.76 62.79 68.89 65.03
MaHDFT Vim-S 74.08 74.40 82.82 86.41 83.93 58.67 50.65 62.39 71.13 65.85
@@
$\bullet$ MSP60K [54] dataset is a newly released pedestrian attribute recognition dataset, which contains 60122 pedestrian images and 57 binary attributes. In the experiment, according to the setting of MSP60K, 30,298 images are used as the training set, 6002 images are used as the validation set, and 23,822 images are used as the test set in the random partition. In the cross-domain partition, 34128 images from five scenes (construction site, market, kitchen, school, ski resort) are used for training, and 24994 images from three scenes (Outdoors1, Outdoors2, Outdoors3) are used for testing.
@@ IV-B Implementation Details
For the detailed parameters, we train the model for 100 epochs using the Adam [58] optimizer. When using VMamba-B, we set the learning rate to 8e-5 and the batch size to 16. When using Vim-S, we set the learning to 2.5e-4 and the batch size to 64. Our source code is implemented using Python and the deep learning framework PyTorch [59]. The experiment were conducted on the server using a GPU RTX3090s. More details can be found in our source code.
IV-C Comparison on Public PAR Benchmarks
@@
TABLE III: Comparison with state-of-the-art methods on RAPv1 and RAPv2 datasets. “V" means Vision branch. “-" means this indicator is not available.